Suppose you travel at an average rate of 30 miles per hour. Which equation would you use to find the time (t) it takes to travel 575 miles?
575t = 30
t = 30/ 575
t/575 = 30
30t = 575

Answer by stanbon(75887)

Suppose you travel at an average rate of 30 miles per hour. Which equation would you use to find the time (t) it takes to travel 575 miles?
-----------------
t = distance/rate
t = 575/30
OR
30t = 575
====================
Cheers,
Stan H.
====================
575t = 30
t = 30/ 575
t/575 = 30
30t = 575